OpenStudy (anonymous):

PLEASE HELP! Solve the following problem. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. In a triangle, the middle length side is 3 more than the shortest side, and the longest side is double the middle length side. The perimeter of the triangle is 45 cm. Find the lengths of the sides.

OpenStudy (anonymous):

Let s be the shortest side and l the longest side and m the middle side. Translate your problem into equations

OpenStudy (anonymous):

Here is one of them s+l +m =45

OpenStudy (anonymous):

m = s+ 3

OpenStudy (anonymous):

l =2 m

OpenStudy (anonymous):

replace in the first equation s+l +m =45 s+ 2m + m=45 s + 3 m =45 s +3(s+3) =45 s + 3 s + 9 =45 4s=36 s=9 Can you find the other sides now?

OpenStudy (anonymous):

m = s+3=9+3=12

OpenStudy (anonymous):

l = 2 m= 24

OpenStudy (anonymous):

Could you list it using V.E.S.T ? @eliassaab

OpenStudy (anonymous):

What is VEST? The first time I hear about it.

OpenStudy (anonymous):

Variable, Equation, Solve, Try it.

OpenStudy (anonymous):

That is what I did. To try it 24+12+9 =45

OpenStudy (anonymous):

24 = 2(12) 12 = 3+9

OpenStudy (nikato):

V Short side=x Middle=3+x Long= 2(3+x)

OpenStudy (nikato):

E Since the perimeter=45, that means all 3 sides =45 so x+3+x+2(x+3)=45 Distribute x+3+x+2x+6=45 Combine like terms 4x+9=45 Subtract 9 on both sides 4x=36 Divide both sides by 4. x=?

OpenStudy (nikato):

Once u find x, plug in what x is into the 3 expression in V. For t, once u find all 3 lengths, add it all up and see if it adds up to 45. That's test it!

OpenStudy (anonymous):

how do i divide 4x and 4 ? wpuld that be x? Then the answer for x would be 9?

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

What is "solve"?

OpenStudy (nikato):

sorry. the website wasnt loading for me at first. u solved when u started distributing the equation. and when u solved it, u got x=9